### v1

@@ Abstract
We study belief polarization among Bayesian agents observing public information about a multidimensional state. [27] show that divergence in the sense of first-order stochastic dominance is impossible for one dimensional beliefs, but we find that in multidimensional settings it can occur for all marginal beliefs, even with infinitely many signals. At the same time, we extend their impossibility result: divergence in the sense of multidimensional stochastic dominance is impossible. For an intermediate stochastic order, polarization may arise only in the short-run. We provide necessary and sufficient conditions on signal structures for persistent polarization and discuss implications for polarization in actions.
@@ 1 Introduction
Polarization is a buzzword in current political discourse. Merriam-Webster named it its 2024 Word of the Year, noting that it is “one idea that both sides of the political spectrum agree on.” [^2] See Merriam-Webster “2024 Word of the Year: Polarization,” accessed June 26, 2025: merriam-webster.com/wordplay/word-of-the-year. [31] clarify that polarization can be defined both as a state—one in which society’s views are concentrated at opposite extremes, or as a process—i.e., the movement towards such a state. Furthermore, polarization can pertain to attitudes, beliefs, preferences etc. This paper focuses on beliefs and on polarization as a process. Specifically, we will use the following notion, based on the definition in [32] (henceforth DW).
Polarization occurs if the beliefs of two agents move further apart and in opposite directions after they observe the same information.
At first glance, this type of polarization appears at odds with rational Bayesian models. A common feature of such models is the prediction that when agents observe the same information their beliefs move closer together. Therefore, widespread polarization seemingly offers evidence of irrationality on at least one side. This intuition is formalized by several theoretical results, notably the impossibility theorem of [27] (henceforth BHK). Considering two Bayesian agents with different priors, they show that observing a common public signal cannot cause beliefs to diverge in the sense of first-order stochastic dominance. [^3] Meaning that if agent $H$’s prior stochastically dominates agent $L$’s prior, it cannot be the case that $H$’s posterior dominates $H$’s prior while $L$’s posterior is dominated by $L$’s prior. Interpreting polarization as irrational has deep implications. It fits the broader view that we are living in a post-truth era, marked by erosion of shared facts and a narrowing scope for constructive public discourse [41, 36].
@@
In BHK, agents’ beliefs are defined over a one-dimensional state space. In practice, political beliefs are multidimensional, spanning foreign policy, the economy, cultural issues, and more. The first contribution of this paper is to show that once beliefs are multidimensional, a strong and practically relevant form of Bayesian polarization becomes possible: agents’ marginal beliefs can diverge simultaneously in the sense of stochastic dominance on each dimension. In other words, the type of divergence that [27] rule out in one dimension may arise coordinatewise across all dimensions.
The analysis also clarifies the limits of Bayesian polarization. Two distinctions matter: whether polarization is measured after a single signal realization (one-shot) or after an infinite sequence of i.i.d. realizations (limit); and which stochastic order is used to compare beliefs. The coordinatewise order described above requires stochastic dominance on each marginal dimension separately, and here both one-shot and limit polarization are possible. The paper’s second contribution is to establish the restrictions that Bayesian updating places on polarization, by showing impossibility results for stronger notions of divergence. The multidimensional stochastic dominance order is the standard generalization of one-dimensional stochastic dominance to higher dimensions. Under this order, both one-shot and limit polarization are impossible—an extension of BHK’s impossibility result to any number of dimensions. The upper-orthant order requires divergence on events involving multiple dimensions being simultaneously high (or low). It is stronger than the coordinatewise order but weaker than multidimensional stochastic dominance. [^4] Both the multidimensional stochastic order and the upper-orthant order have been suggested as generalizations of univariate stochastic dominance [48]. The former preserves the property that one distribution dominates another if and only if it dominates it in expectation for every increasing function. The latter is weaker, and preserves the property that one distribution dominates the other if and only if its distribution function dominates the other’s at every point. Under this intermediate criterion, one-shot polarization is possible but limit polarization is not. Table 1 summarizes.
Taken together, the results show that while there are forces pushing Bayesian agents towards agreement, these still allow for significant polarization in the face of common information. We demonstrate this idea from a different angle by considering polarization in beliefs on individual states. In any finite state space, there are always two states—the minimum- and maximum-likelihood states—on which both agents’ beliefs must move in the same direction. On all other states, however, their beliefs may move in opposite directions.
@@
Mathematically, the BHK impossibility result generalizes, but substantively it is the coordinatewise order—not multidimensional stochastic dominance—that best reflects how polarization is measured in practice. Survey evidence often tracks divergence across separate issue dimensions such as climate change, trade, or cultural issues (see, e.g., [43]). The possibility result shows that this form of polarization is fully consistent with Bayesian updating, even in the long run. By contrast, the stronger orders capture divergence on combinations of issues, e.g., whether the state is simultaneously high on both economic and cultural dimensions. The impossibility results imply that agreement may persist on such joint events, even when marginals diverge. Evidence of polarization on such issue combinations would be harder to reconcile with rational updating.
@@
The third contribution is a complete characterization of the signals that can lead to limit coordinatewise polarization in two-dimensional state spaces. An immediate necessary condition is partial identification: there must be at least one additional state that is observationally equivalent to the true state of the world. Otherwise, beliefs converge to the truth, precluding polarization. [^5] Formally, if the signal is $X$, the state space is $\Theta$, and the true state is $\hat{\theta}$, then there must be at least one state $\theta\in\Theta\setminus\{\hat{\theta}\}$ such that for every realization $x$: $\Pr[X=x|\theta]=\Pr[X=x|\hat{\theta}$]. Otherwise, Doob’s Consistency Theorem ([33]) implies that, from any full-support prior, beliefs converge to a point mass on $\hat{\theta}$. Fixing the true state, any polarizing signal must therefore induce a set of states observationally equivalent to it, which we call the identified set. The problem of identifying polarizing signals thus reduces to characterizing the identified sets that permit polarization.
Under a mild technical condition, Theorem 5 shows that an identified set can lead to polarization if and only if it satisfies three geometric conditions. Both the set and its complement must be spanning—each must include states that attain the minimal and maximal values on every dimension. The set must be balanced—it cannot include all states that lie strictly below some point while excluding all of the states that lie weakly above it, or vice versa. And it must be non-compensatory—it cannot be the case that whenever states in the set are high on one dimension they are low on the other.
To illustrate the theorem, consider a simple $2\times 2$ state space $\Theta=\{x_{1},x_{2}\}\times\{y_{1},y_{2}\}$ with $x_{1}<x_{2}$ and $y_{1}<y_{2}$. Think of the $x$ dimension as describing the state of the economy, and the $y$ dimension the state of foreign policy. Suppose Democrats view both as low and Republicans view both as high. Two agents, $L$ and $H$, have full support priors over this state space. Agent $L$’s views are closer to the Democratic position: her prior is stochastically dominated by agent $H$’s on each dimension.
As shown in Figure 1, only two sets are spanning with spanning complements. These are the diagonal $\{(x_{1},y_{1}),(x_{2},y_{2})\}$ and the off-diagonal $\{(x_{1},y_{2}),(x_{2},y_{1})\}$. The off-diagonal violates the non-compensatory condition, since whenever a state is low on the economic dimension it is high on the foreign policy dimension, and vice-versa. The diagonal, on the other hand, satisfies all three conditions. Thus, a corollary of Theorem 5 is that in $2\times 2$ state spaces, limit coordinatewise polarization can arise if and only if the identified set is the diagonal. That is, polarization is possible only when the signal reveals that the two dimensions are perfectly correlated (and nothing else).
[…]
(a) set and its complement not spanning
[…]
(b) compensatory
[…]
(c) biased; complement not spanning
[…]
(d) uniquely satisfies the three conditions
Figure 1: Applying Theorem 5 to the $2\times 2$ space.
The result connects to a central stylized fact: the steady rise in polarization among US political elites since the 1970s has coincided with rising ideological consistency, with positions on one issue increasingly predicting positions on others [40]. The theorem’s conditions permit arbitrarily high correlation between the dimensions but the non-compensatory condition rules out strong negative correlation. [^6] In the $2\times 2$ case limit polarization can arise only when agents learn that the two dimensions are perfectly correlated. Generally, learning that the two dimensions are perfectly correlated can always lead to polarization while learning that they are perfectly negatively correlated cannot. This points to a complementary Bayesian mechanism to existing accounts of polarization: if elites encountered evidence of strong correlations among issues, their posteriors would have polarized while becoming more internally consistent, in line with the observed pattern. A suggestive illustration is the fall of ‘Third Way’ politics. This movement, championed in the 1990s by Bill Clinton and Tony Blair, combined conservative economic policies with progressive social stances. Its decline by the early 2000s [37] may have signaled the instability of such mixed positions, strengthening ideological consistency and thereby contributing to polarization.
The spanning and balance conditions imply that polarizing signals must (i) preserve the possibility that the state takes very low or very high values on each dimension, and (ii) avoid decisively indicating that the state is simultaneously high (or low) on both dimensions. This echoes journalistic norms of ‘balanced coverage’, which may seem counterintuitive, since media bias is commonly viewed as a key driver of polarization. Yet the biased-media account typically works through selective exposure, with each side consuming different evidence that supports its prior (see, e.g., [39]). By contrast, we study polarization in the presence of public information. In this setting, evidence that decisively supports one side cannot lead to polarization, but balanced evidence can. Historical accounts suggest that balanced coverage has, at times, prolonged disagreement. For instance, [45] document how tobacco industry leaders exploited the media’s adherence to the Fairness Doctrine to sustain controversy about the health risks of smoking, despite overwhelming scientific evidence.
While the analysis centers on beliefs, polarization is often discussed in terms of actions, such as voting, campaign donations, or social media posts. In the concluding section, we develop a simple political-economy model that connects the two. The key insight is that whether action polarization arises depends on the separability of preferences: robust action polarization arises when utility functions are additively separable across dimensions, but not when they display strong complementarities.
The paper proceeds as follows. Section 2 reviews related literature. Section 3 presents the model and the definitions of polarization, and Section 4 reviews preliminaries on stochastic orders and Bayesian learning. Section 5 develops the possibility and impossibility results, and Section 6 characterizes the identified sets that lead to limit coordinatewise polarization. Section 7 briefly discusses the probability of belief polarization, and polarization in actions.
@@ 2 Related Literature
Our notion of polarization builds on DW, who study a one-dimensional state space and define polarization as divergence in the stochastic dominance order after a single signal realization. DW show that polarization is impossible under Bayesian updating if signals satisfy the monotone likelihood ratio property (MLRP), and in particular in the standard linear-normal model. They also present examples suggesting that polarization may arise when MLRP fails, though, as later noted by BHK these examples display only divergence of expectations, not polarization in the stochastic dominance order. BHK adopt the DW definition and prove that polarization is impossible when agents are Bayesian and the state space is finite and one dimensional. [^7] The examples that DW present in detail do not display polarization according to their definition. They also sketch additional examples with continuum state spaces for which they do not specify the priors. Technically, these do not violate BHK, which assumes a discrete state space. Whether polarization is possible in continuum state spaces is an open question.
@@
BHK’s impossibility result cemented the intuitive notion that belief polarization is inconsistent with Bayesian learning. This notion is in line with classic results in [33] and [29], which predict different forms of belief convergence among Bayesians learning from the same information. To reconcile the perceived theoretical impossibility with the apparent prevalence of polarization, several papers have introduced learning biases that generate polarization. These include ambiguity aversion (BHK), confirmation bias [46], misperceptions about selective sharing on a network [30], and correlation neglect [38].
Other work has challenged this view by showing that some forms of polarization are consistent with Bayesian updating. The closest to ours are [26] and [28], who demonstrate polarization in expectations on one dimension when agents hold beliefs over a two-dimensional state. [^8] Both focus on $2\times 2$ state spaces where the choice of stochastic order is irrelevant, but [26] note that their observation can be generalized to polarization in expectations.
[35] provide examples of Bayesian networks that may lead to polarization regarding the probability of a hypothesis. Put in our terms, this is another example of one-dimensional polarization in a two-dimensional state space. [47] show the possibility of belief dilation: for a family of prior beliefs, observing evidence may lead the upper and lower bounds for the probability of a given event to diverge. [25] study asymptotic disagreement rather than polarization. They focus on the magnitude of disagreement rather than its direction, showing that vanishing uncertainty about likelihood functions can lead to substantial asymptotic disagreement. [44] develop a model of persistent polarization due to rational inattention. [34] presents a philosophical account of predictable polarization, where rational agents’ beliefs diverge in directions they can anticipate ex-ante.
This paper generalizes earlier results on the possibility of Bayesian polarization by demonstrating the possibility of polarization in first order stochastic dominance on all dimensions simultaneously, for any number of dimensions, and following an infinite sequence of i.i.d. signals. It contributes to the debate on the rationality of belief polarization by showing both the possibility of Bayesian polarization and its limits. While most previous papers focused on short-term polarization, we distinguish between one-shot and asymptotic polarization and show that this distinction matters: under the upper orthant order, one-shot polarization is possible but asymptotic polarization is not. Finally, we characterize the signals that can lead to limit polarization and connect polarization in beliefs to polarization in actions.999[28] characterize the signals that lead to one-shot polarization in their setting, motivated by experimental evidence and focused on polarization in populations. Since their setting is more specific and the notions of polarization differ, the conditions are not directly comparable.
3 Model

### v2

@@ Abstract
Discussions of political disagreement emphasize two patterns: polarization, where beliefs diverge toward opposite extremes on each issue dimension; and issue alignment, where individuals’ views across issues become more internally consistent. We show that both can simultaneously arise under Bayesian learning from public information. We characterize the public signals that can induce persistent polarization on all dimensions and find that evidence of issue alignment can polarize Bayesian agents. However, we show that even stronger notions of polarization—requiring divergence beyond marginal beliefs—are inconsistent with Bayesian rationality. Whether multidimensional belief polarization translates into divergent aggregate positions depends on cross-issue separability.
@@ 1 Introduction
Polarization is widely viewed as a defining feature of contemporary politics. Merriam–Webster named polarization its 2024 Word of the Year, noting that it is “one idea that both sides of the political spectrum agree on.” Consistent with this sentiment, a [35] survey reports that 80% of Americans believe that Republican and Democratic voters disagree not only over policies but also over basic facts.
The term polarization is often used as an umbrella for several related phenomena. A central notion is divergence: policy preferences or ideological positions moving farther apart toward opposite extremes. A distinct but related notion is issue alignment: individuals’ beliefs becoming more internally consistent, e.g., those who are liberal on one policy dimension are more likely to be liberal on others. Throughout the paper, we reserve the term polarization to refer specifically to divergence.
Focusing on political elites in the United States, [27] documents two stylized facts: elites—politicians at all levels, judges, and media figures—have become increasingly polarized since the 1970s, and this rise in polarization has coincided with increased issue alignment. For example, the correlation between legislators’ positions on economic issues and civil rights in the 1960s was substantially weaker than the correlation between positions on economic policy and affirmative action in the 1990s. Survey evidence points to similar patterns among the public. A [15] study of attitudes on 24 key issues finds that partisan gaps have either persisted or widened across all dimensions, with opinions often moving in opposite directions [[, see also]]Pew2014Polarization,pew2017partisan. However, while there is broad agreement that strong partisans display polarization and issue alignment patterns similar to those of political elites, scholars debate whether the patterns observed in the general public reflect genuine belief change or only increased alignment between individuals’ views and their partisan identities.222[1] argue that mass polarization is substantial, whereas [13] argue that it is largely a myth. See [18] for a review.
A popular explanation for these trends emphasizes selective exposure to information. Social media platforms are said to facilitate “echo chambers,” in which exposure to like-minded others reinforces prior beliefs [41, 20]. But recent studies challenge this account: the extent of selective exposure online is limited [16], and some experimental studies find no link between exposure to online echo chambers and polarization [32]. Consistent with these findings, the [35] survey reports that 67% of Americans attribute partisan disagreement over basic facts primarily to differing interpretations of the same information.
Prima facie, this type of polarization appears irrational. Intuition as well as classic theoretical results suggest that rational Bayesian learning from the same information leads to belief convergence, so polarization should not arise. [4] (henceforth BHK) present an impossibility result that formalizes this intuition. Considering two Bayesian agents with different priors, they show that observing a common public signal cannot cause beliefs to diverge in the stochastic dominance order. [^3] Throughout the paper, stochastic dominance refers to first-order stochastic dominance; see formal definition below.
@@
In BHK, agents’ beliefs are defined over a one-dimensional state space. In practice, political beliefs are multidimensional, spanning foreign policy, the economy, cultural issues, and more. It may seem that if one-dimensional Bayesian polarization is impossible, then so is simultaneous polarization on multiple dimensions; in particular, patterns like those reported in the [15] study are inconsistent with Bayesian updating from public information. Our first contribution is to overturn this intuition: when Bayesian agents’ beliefs are defined over a multidimensional state space, a public signal can lead all marginal beliefs to simultaneously diverge in the sense of stochastic dominance. Moreover, this polarization can persist even if agents observe an infinite sequence of i.i.d. signals. Thus, the type of polarization that BHK rule out in one dimension after any public signal realization may arise coordinatewise across all dimensions, even in limit beliefs.
Having established that multidimensional Bayesian polarization is possible, we turn to two natural questions: what type of information generates it, and what are its limits? Our second contribution is a characterization of the public signals that can generate persistent polarization of all marginal beliefs. The result suggests a novel informational mechanism consistent with the stylized facts: evidence of issue alignment can lead to polarization. Our third contribution is to delineate the limits of Bayesian polarization by establishing possibility and impossibility results for stronger notions of belief divergence.
We establish these results within a simple model. Two Bayesian agents, $L$ and $H$, have prior beliefs over a finite multidimensional state space. There is a meaningful order across all dimensions. For example, the state may represent an optimal policy bundle, with dimensions representing economic, foreign, and climate policy, each ordered from the most liberal policy to the most conservative. Agent $L$ believes that the state is lower than agent $H$ on all dimensions, i.e., $L$ initially tends toward liberal policies. We will say that polarization occurs if there is a public signal that moves $L$’s beliefs further down on all dimensions and $H$’s beliefs further up. Building on [10], who study polarization in a one-dimensional state space, we use stochastic dominance to compare marginal beliefs. So our notion of polarization requires that agent $L$’s posterior marginals are stochastically dominated by agent $L$’s prior marginals and vice versa for $H$.
Theorem 1 shows that this form of polarization is possible. The following example illustrates the result.
Example 1.
Consider a $2\times 2$ state space $\Theta=\{x_{1},x_{2}\}\times\{y_{1},y_{2}\}$ with $x_{1}<x_{2}$ and $y_{1}<y_{2}$. For concreteness, interpret the $x$-dimension as economic policy and the $y$-dimension as immigration policy, with lower values corresponding to more liberal positions. Let $P^{L}$ and $P^{H}$ denote the prior beliefs of agents $L$ and $H$ over the optimal policy bundle, shown in Figure 1. Agent $L$ assigns more mass to low values on each dimension than agent $H$, so that every marginal of $P^{L}$ is stochastically dominated by the corresponding marginal of $P^{H}$.
[…]
Figure 1: Priors for Example 1.
Consider a public signal realization that reveals that the state lies on the “diagonal”: $\{(x_{1},y_{1}),(x_{2},y_{2})\}$, and is equally likely to be generated in either of these states. That is, the signal implies perfect positive association between the two dimensions while remaining uninformative about the value on each dimension. Let $Q^{L}$ and $Q^{H}$ denote the posterior beliefs obtained by Bayesian updating. These are shown in Figure 2.
[…]
Figure 2: Posteriors for Example 1.
Despite observing the same signal, the agents’ marginal beliefs diverge on each dimension. Writing $P_{i}^{L},P_{i}^{H}$ and $Q_{i}^{L},Q_{i}^{H}$ for the marginal distributions on dimension $i=1,2$, we have (here $\succeq_{st}$ is the stochastic dominance order)
$Q^{L}_{i}=(3/4,1/4)\prec_{st}P^{L}_{i}=(5/8,3/8)\prec_{st}P^{H}_{i}=(3/8,5/8)\prec_{st}Q^{H}_{i}=(1/4,3/4).$
Thus, agent $L$ becomes more confident that the state is low on each dimension, while agent $H$ becomes more confident that it is high, even though both update rationally from the same information.
Example 1—which extends to any finite number of dimensions—shows that coordinatewise polarization is possible after a single public signal. As mentioned above, we also establish the possibility of limit coordinatewise polarization—polarization in limit beliefs for an infinite sequence of i.i.d. signals.
Turning to the information that generates polarization, Theorem 2 presents a complete characterization of the signals that can lead to limit coordinatewise polarization in two-dimensional state spaces. An immediate necessary condition is partial identification: there must be at least one additional state that is observationally equivalent to the true state of the world. Otherwise, beliefs converge to the truth, precluding polarization. [^4] Formally, if the signal is $X$, the state space is $\Theta$, and the true state is $\hat{\theta}$, then there must be at least one state $\theta\in\Theta\setminus\{\hat{\theta}\}$ such that for every realization $x$: $\Pr[X=x|\theta]=\Pr[X=x|\hat{\theta}$]. Otherwise, Doob’s Consistency Theorem [11] implies that, from any full-support prior, beliefs converge to a point mass on $\hat{\theta}$. Therefore, any polarizing signal must induce a set of states that are observationally equivalent to the true state. We call this the identified set. The problem of characterizing limit polarizing signals reduces to characterizing the identified sets that permit polarization.
Under a mild technical condition, Theorem 2 establishes that an identified set can lead to coordinatewise polarization if and only if it satisfies three geometric conditions. Two of these—spanning and balance—govern how much the signal resolves disagreement about the value of the state. Spanning requires that both the identified set and its complement include states attaining the minimal and maximal values on each dimension, so that extreme outcomes remain possible in every direction. Balance rules out signals that decisively point toward uniformly high or uniformly low states. Taken together, these conditions ensure that the signal preserves uncertainty about the magnitude of the state, both within and across dimensions.
The third condition, non-compensation, governs how the signal links the dimensions to one another. It rules out signals that systematically trade off one issue against another by pairing high values on one dimension with low values on the other. This condition is closely related to issue alignment: signals that imply perfect positive association between dimensions satisfy it, whereas signals implying perfect negative association violate it. Intuitively, polarization requires information that links the different issues without resolving disagreement about their values.
In the $2\times 2$ state space, the theorem delivers a sharp implication. The diagonal $\{(x_{1},y_{1}),(x_{2},y_{2})\}$ is the unique identified set that satisfies all three conditions. Consequently, the signal in Example 1 is essentially the only public signal that can generate persistent polarization on both dimensions. In this case, polarization necessarily coincides with increased issue alignment. In larger state spaces, increased alignment is no longer necessary for polarization; yet there always exist signals that generate both polarization and perfect alignment.
The result points to a mechanism through which public information about issue alignment can generate polarization. Information that reveals that dimensions are aligned—without revealing whether the state is overall high or low—can push beliefs toward opposite extremes. This information could be interpreted as a growing public understanding of an ideological structure that deems mixed positions—for example, liberal on economics and conservative on immigration—inconsistent [9]. In line with the idea that agents can learn about alignment, [24] show that such ideologically structured positions are positively associated with political sophistication. Regardless of the source of issue alignment, our contribution is to show that once it emerges, it can generate polarization through simple Bayesian updating.
Up to this point, we have considered polarization of all marginal beliefs, which we refer to as polarization in the coordinatewise order. To clarify the constraints that Bayesian updating imposes on polarization, we next consider polarization in two stronger stochastic orders commonly used to compare multivariate distributions. Theorem 3 shows that polarization in the upper orthant order is possible in the short-run but not in the long-run. Theorem 4 shows that polarization in the stochastic order is impossible following any single signal realization. Table 1 summarizes these results. Here, we distinguish one-shot polarization, which occurs after a single signal realization, from limit polarization, which arises following an infinite sequence of i.i.d. signals.
Taken together, the three orders define a feasibility frontier for Bayesian polarization. One-shot polarization is possible for any order weaker than the upper orthant order, but impossible for any order stronger than stochastic dominance. Limit polarization is possible for orders weaker than the coordinatewise order, but impossible for orders stronger than the upper orthant order.
@@
Mathematically, our result for the stochastic dominance order shows that the BHK impossibility result extends to higher dimensions. Substantively, however, it is the coordinatewise order—not stochastic dominance—that best reflects how polarization is measured in practice. Empirical work typically tracks divergence along individual issue dimensions such as climate policy, trade, or cultural issues. The coordinatewise possibility result shows that this form of polarization is fully consistent with Bayesian updating, even in the long-run limit. By contrast, the stronger orders capture divergence on combinations of issues, e.g., whether the state is simultaneously high on several dimensions. The impossibility results imply that agreement may persist on such joint events, even when marginal beliefs diverge. Evidence of polarization on such issue combinations would be harder to reconcile with rational updating.
We conclude by studying when multidimensional belief polarization is reflected in one-dimensional aggregate positions. We show that the scope for a robust notion of polarization in positions depends on the separability of the aggregation rule across dimensions: additively separable aggregators allow persistent polarization, multiplicatively separable aggregators allow it only in the short run, and sufficiently strong cross-dimensional complementarities preclude it.
@@ 2 Related Literature
Our main notion of polarization builds on that of [10], who study Bayesian agents with beliefs over a one-dimensional state space and define polarization as divergence in the stochastic dominance order following a single public signal realization. They show that such polarization is impossible when signals satisfy the monotone likelihood ratio property (MLRP). BHK adopt this definition and prove that polarization is impossible for any signal realization when agents are Bayesian and the state space is finite and one-dimensional.555[10] present examples that violate MLRP, which they argue generate polarization. As noted by BHK, however, these examples exhibit divergence in expectations rather than divergence in stochastic dominance. [10] also sketch examples of polarization in continuum state spaces; since BHK’s impossibility result applies only to finite state spaces, the possibility of polarization in the continuum remains open.
@@
BHK’s impossibility result reinforced the view that polarization is inconsistent with Bayesian learning. This aligns with classic results by [11] and [7], which establish distinct forms of belief convergence among Bayesians observing the same information. To reconcile the perceived theoretical impossibility with the apparent prevalence of polarization, a growing literature has introduced learning biases or other departures from standard Bayesian updating. [^6] This literature is often motivated by same-evidence polarization experiments, starting with [23]; see [36] and [6] for a review and discussion. These include ambiguity aversion (BHK), confirmation bias [36, 14], misperceptions about selective sharing on networks [8], short-term memory [21], and correlation neglect [19, 22, 34].
Other work has challenged this view by showing that some forms of polarization are consistent with Bayesian updating. In particular, several papers demonstrate that polarization along a single dimension may arise when agents hold beliefs over a multidimensional state space. [3] provide a simple theoretical example and experimental evidence, while [17] present additional examples using Bayesian networks. A more general analysis is provided by [6], who characterize the public signals that can generate such polarization. Whereas they focus on beliefs after a single realization, [2] consider agents observing an infinite sequence of i.i.d. signals and establish that limit beliefs can diverge when agents are uncertain about signal likelihoods. In our terms, they demonstrate limit polarization on one dimension in a two-dimensional state space. However, their focus is on the magnitude of disagreement rather than its direction, showing that even vanishing uncertainty about likelihood functions can lead to substantial asymptotic disagreement. Taken together, these papers show that Bayesian learning does not preclude polarization, but suggest that for agents to polarize they need to be uncertain about some ancillary dimension, distinct from the dimension of disagreement.
We show that this restriction is unnecessary: Bayesian agents may simultaneously polarize on all dimensions of the state space. Moreover, on each dimension polarization occurs in the strong sense of stochastic dominance, and can persist in limit beliefs. [^7] Most previous papers demonstrated polarization over dimensions with only two elements, where the choice of stochastic order is irrelevant. In the short run, we find that an even stronger form of polarization—in the upper orthant order—is achievable.
Several papers have argued for the rationality of polarization outside the standard Bayesian paradigm or using different notions of polarization. [38] characterize belief dilation: the divergence of probability bounds for a family of beliefs following any signal realization. [31] obtain persistent polarization due to rational inattention. [30] consider polarization of Bayesian beliefs on individual events and in the total variation distance. [12] argues that it can be rational for beliefs to diverge in ex-ante predictable directions.
Overall, existing work has reached seemingly conflicting conclusions on the compatibility of polarization with Bayesian learning, largely due to differing definitions. We contribute by clarifying which types of polarization are feasible for Bayesian agents with beliefs over a finite multidimensional product space. Furthermore, while most previous papers focus either on short-run or limit polarization, we distinguish between the two and show that the distinction has bite: under the upper orthant order, polarization is possible only in the short run.
Finally, we analyze which public information can generate limit coordinatewise polarization, and show how evidence of issue alignment can itself be polarizing. [6] also characterize polarizing signals, but consider a more specialized setting and a weaker notion of polarization. [^8] In our terms, [6] characterize the signals that lead to one-shot polarization on one dimension in a $2\times 2$ space, whereas we characterize the signals that lead to limit coordinatewise polarization in arbitrary finite two-dimensional spaces. Their characterization serves as a stepping stone toward an analysis of population polarization, which we do not consider. [40] and [42] present different models in which a form of rational learning generates both issue alignment and polarization: the former models agents learning about a common state from each other’s cheap-talk messages, while the latter considers rationally inattentive agents learning about their idiosyncratic ideological positions. In contrast to these papers, we do not model why issue alignment arises. Instead, we observe that when it does, it can generate polarization through a Bayesian mechanism.
3 Preliminaries
3.1 One-shot and Limit Polarization